Angeline wishes to make six annual deposits in the amounts of $1,000, $2,000, $3,000,
$4,000, $5,000 and $6,000, in that order, into a fund that pays interest at rate of 5% per
annum. Determine the value of amount today if
i) the interest is compounded annually
ii) the interest is compounded quarterly.
1
Expert's answer
2020-10-15T03:00:56-0400

solution


Part i)


Compounding per annum


Rate "r=0.05"

Present value "PV = \sum p_i(1+r)^{-i}"



"= \frac{1000}{(1.05)^1} + \frac{2000}{(1.05)^2} + \frac{3000}{(1.05)^3}""+ \frac{4000}{(1.05)^4} + \frac{5000}{(1.05)^5} + \frac{6000}{(1.05)^6}"


"= 17044.74"

answer: the present value is $ 17,044.74


Part ii)


Compounding quarterly


"PV = \sum p_i(1+\frac{r}{4})^{-i*4}"


"= \frac{1000}{(1.0125)^4} + \frac{2000}{(1.0125)^8} + \frac{3000}{(1.0125)^{12}}""+ \frac{4000}{(1.0125)^{16}} + \frac{5000}{(1.0125)^{20}} + \frac{6000}{(1.0125)^{24}}"

"=26116.03"


answer: the present value is $ 26,116.03